# Real Noncommutative Convexity II: Extremality and nc convex functions

### v1

@@ Abstract.
We continue with the theory of real noncommutative (nc) convexity, following the recent and profound complex case developed by Davidson and Kennedy. The present paper focuses on the theory of nc extreme (and pure and maximal) points and the nc Choquet boundary in the real case, and on the theory of real nc convex and semicontinuous functions and real nc convex envelopes. Our main emphasis is on how these interact with complexification. For example some of our paper analyzes carefully how various notions of ‘extreme’ or ‘maximal’ interact with our earlier concept of the complexification of a convex set. Several new features appear in the real case in later sections of our paper, including the novel notion of the complexification of a nc convex function, and the complexification of the convex envelope of a nc function. With an Appendix by T. Russell.
@@ 1. Introduction
Very recently in [8] we developed the real case of the theory of real noncommutative (nc) convex sets in the sense of Davidson and Kennedy. We continue here with the theory of real noncommutative (nc) convexity, following the recent and profound complex case developed by those authors in the memoir [14]. As before, one of our primary goals is to develop the role and properties of the complexification in this context. One main motivation for our work, as we mentioned in more detail in [8], is that ‘classical convexity’ is in some ways essentially a real theory, and so the real variant of the theory is likely to be important in the future in operator algebras, mathematical physics, and ‘quantum analysis’ (see e.g. [16] as a representative example of what part of this might look like). The present paper focuses on the real case of the topics from the middle sections of the deep memoir [14], namely on the theory of nc extreme (and pure and maximal) points and the nc Choquet boundary in the real case, and on the theory of real nc convex and semicontinuous functions and real nc convex envelopes; especially on how these interact with complexification. One use of this is as a tool to transfer results back and forth between the real and complex nc convexity theory, which often allows one to avoid difficulties. Some of our paper analyzes carefully how various notions of ‘extreme’ interact with our earlier concept of the complexification of a convex set. More sophisticated results, such as more advanced nc Choquet theory, will be discussed in a future paper. Several new features appear in the real case in later sections of our paper, including the novel notion of the complexification of a nc convex function on a nc convex set $K$, and the complexification of the convex envelope of a nc function on $K$. As just one example, we show that the convex envelope construction commutes with complexification. The convex envelope of a function $f$ defined on a convex set is of utmost importance in classical convexity theory, where it is sometimes called the ‘lower envelope’. (See e.g. [30]. Of course the upper/concave envelope may be treated via the convex envelope of $-f$.) More generally we continue giving many foundational structural results for real nc convex sets and the functions on them.
@@
There is of course a large and highly impressive literature (going back thirty years or so) on noncommutative or quantum versions of extreme points, Krein-Milman, etc, and also on noncommutative versions of convex functions. As just a few recent examples see for instance the remarkable papers [20, 22, 35, 34] (and the list of papers referred to on p. 2 of [8]), and references therein. (At the time of submission of the present paper Scott McCullough also pointed out the very very recent [21].) Indeed there are even classical precursors e.g. in the operator (or matrix) convex and monotone functions going back to Löwner. Moreover this large existing theory of ‘operator/matrix convexity’ of which we speak often focuses on real examples (e.g. spectrahedra), and sometimes is even completely cast in the real framework, using ‘real’ tools. However we believe that there is no technical overlap of our work with this literature, since our goal here is mainly to extend the perspective of [14] to the real case, particularly emphasizing and developing the role of the complexification in this endeavor. (In some stretches of Section 2 where the complexification plays little part we are essentially just verifying and resolving potential complications in, and otherwise expositing, the real case of [14, Section 6]. Our contributions in this part of Section 2 are not intended to be novel, but rather to be useful and systematic for future ‘real users’.) Indeed aspects of the work earlier cited in this paragraph will hopefully interact very profitably with the matter of the present paper(s) in the future, and in any case this cited work certainly furnishes very many interesting and important real examples. Much of this cited work is stated in the Wittstock-Effros-Webster-Winkler language [40] of matrix convexity. We refer readers to [14] for discussion some of the differences between nc and matrix convexity.
@@
Turning to the structure of our paper, in Section 2 we begin by characterizing maximal elements in a complexification. This gives a new route to the construction of the $C^{*}$-envelope of a real operator system (or unital operator space). We also discuss the real case of the theory of Davidson and Kennedy’s nc convexity variant of extreme, pure, and maximal points (which when ‘irreducible’ are essentially Arveson’s boundary representations). And of the associated nc Choquet and Shilov boundary, and its application to the nc Krein-Milman theorem and its converse. In most of the rest of this section we give a few points of clarification about the real case of the arguments, and note a few places where the arguments are different in the real case. In Section 3 we examine the connections between extreme, pure, and maximal points in the real and complex case, that is, how the these notions interact with the complexification of a nc convex set. One may view some of our work in this section as complementary to parts of the very recent paper [33]. In that paper some analogous results are established for $C^{*}$-extreme maps on real $C^{*}$-algebras. Every pure (and in particular any nc-extreme) ucp map on a $C^{*}$-algebra is $C^{*}$-extreme, but not vice versa. They do not treat general operator systems, thus their results would not apply to points in general nc convex sets. In Sections 4 and 5 we consider the real case of Davidson and Kennedy’s nc convex functions, including their multivalued nc convex functions, and convex envelopes.
We give many foundational structural results here, for example developing the theory of the complexification in these settings and elucidating how it interacts with the basic objects and constructions. E.g. we show that the convex envelope construction commutes with complexification. As in [8], the complexification then allows us to derive the real variant of profound and extremely technical results from [14, Section 7] very quickly from the complex case.
@@
We write $M_{n}(\mathbb{R})$ for the real $n\times n$ matrices, or sometimes simply $M_{n}$ when the context is clear. Similarly in the complex case. We sometimes use the quaternions $\mathbb{H}$ as an example: this is simultaneously a real operator system, a real Hilbert space, and a real C-algebra, usually thought of as a real *-subalgebra of $M_{4}(\mathbb{R})$ or $M_{2}(\mathbb{C})$. Its complexification is $M_{2}(\mathbb{C})$. The letters $H,L$ are usually reserved for real or complex Hilbert spaces. Every complex Hilbert space $H$ is a real Hilbert space, i.e. we forget the complex structure. More generally we write $X_{r}$ for a complex Banach space regarded as a real Banach space. We write $X_{\rm sa}$ for the selfadjoint elements in a *-vector space $X$. In the complex case $M_{n}(X)_{\rm sa}\cong(M_{n})_{\rm sa}\otimes X_{\rm sa}$, but this fails for real spaces. A subspace of $B(H)$ is unital if contains the identity, and a map $T$ is unital if $T(1)=1$. Our identities 111 always have norm 111. For $z\in M_{n}(\mathbb{C})$ we write $\text{Re }z$ for $x\in M_{n}(\mathbb{R})$ where $z=x+iy$ for $y\in M_{n}(\mathbb{R})$. Finally, for a cardinal $n$ we define the isometry $u_{n}=\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$ where $1_{n}$ is the $n$-dimensional identity operator. We sometimes also write this as $u$.
A concrete complex (resp. real) operator system $V$ is a unital selfadjoint subspace of $B(H)$ for $H$ a complex (resp. real) Hilbert space. For $n\in\mathbb{N}$ we have the identification $M_{n}(B(H))\cong B(H^{(n)})$ where $H^{(n)}$ is the $n$-fold direct sum of $H$. From this identification, $M_{n}(V)$ inherits a norm and positive cone. The latter is the set $M_{n}(V)^{+}\coloneq\{x\in M_{n}(V):x=x^{*}\geq 0$ in $B(H^{(n)})\}$. For $n\in\mathbb{N}$ we define the amplification of a linear map $\varphi:V\to W$ by
@@ 2. Extreme, pure, maximal points, and the nc Choquet boundary
Viewing $K=\mathrm{ncS}(V)$ for an appropriate real operator system $V$ as in [8, Section 4], this corresponds to $\psi\leq\varphi$ for a cp map $\psi$, implying that $\psi$ is a positive scalar multiple of $\varphi$. Indeed it is obvious that such a $\varphi$ is pure in the first sense above (note $x=t_{i}\alpha_{i}^{*}x_{i}\alpha_{i}$ evaluated at 1 implies $1=t_{i}\alpha_{i}^{*}\alpha_{i}$, so $t\alpha_{i}$ is a multiple of an isometry). Conversely, if $\varphi=\varphi_{1}+\varphi_{2}$ for cp maps $\varphi_{i}$, use [9, Lemma 6.3] to write $\varphi=b_{1}\psi_{1}b_{1}+b_{2}\psi_{2}b_{2}$ for ucp maps $\psi_{i}$ and matrices $b_{i}\geq 0$ with $b_{1}^{2}+b_{2}^{2}=1$. If $\varphi$ is pure in the first sense then $\varphi_{i}=b_{i}\psi_{i}b_{i}$ is a positive multiple of $\varphi$. For most practical purposes this allows us in most situations to drop the requirement that $\alpha_{i}$ is a positive scalar multiple of an isometry above, instead saying that $\alpha_{i}^{\mathsf{T}}x_{i}\alpha_{i}$ is a positive scalar multiple of $x$. See Remark 6.1.6 in [14].
@@
[…]
Suppose that $\varphi:V\to B(H)$ is a unital complete order embedding on a real operator system $V$. The real $C^{*}$-algebra generated by the image of a maximal dilation $j:V\to B(H)$ of $\varphi$ is (a copy of) the important $C^{*}$-envelope $C^{*}_{\rm min}(V)$, by the well known arguments for this from [18, 3]. That is, it has the universal property that for any unital complete order embedding $\kappa:V\to B(L)$ there is a unique *-homomorphism $\pi:C^{*}(\kappa(V))\to C^{*}(j(V))$ with $\pi\circ\kappa=j$. The arguments (due to Arveson-Dritschel-McCullough and others, see e.g. the historical notes to [7, Section 4.1]) for the next result work the same in both the real and complex case. As explained and proved in several places, e.g. in [3, Proposition 2.2] or [7, 4.1.11-4.1.13], we have:
@@
We next prove a useful characterization of the maximal elements of a complexification. We recall the fact from the first three lines of the last paragraph of the proof of [8, Theorem 3.4] that an element $x\in K_{2n}$ of a real nc convex set $K$ is of the form $c_{n}(a,b)$ for $a,b\in K_{n}$ if and only if for the unitary matrix $W=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$ we have $W^{*}xW=x$. In this case, $c(u^{*}xu)=x$. Real nc affine maps preserve such elements. Indeed if $\omega$ is nc affine then $W^{*}\omega(c(x))W=\omega(W^{*}c(x)W)=\omega(c(x))$. Hence also $c(u^{*}\omega(c(x))u)=\omega(c(x))$. We use this in the proof below.
@@
A direct proof of this result is possible (see e.g. the proof of Lemma 3.3). $x\oplus x$ is real maximal in $K$. However it is a pleasant exercise that this is equivalent to $x$ being maximal. ∎
@@
This gives a (new) route to the construction of the $C^{*}$-envelope of a real operator system (or unital operator space). Namely the real case of the Dritschel-McCullough-Arveson route, but using the ‘real arguments’ above.
@@
The discussion in 5.1.2 in [14] is the same in the real case. Turning to Section 6 of [14], Remark 6.1.3 in [14] and the following propositions have the same proof as in the complex case (see Proposition 6.1.46.1.46.1.4 and 6.1.56.1.56.1.5 in [14]).
@@
(1) This is essentially [2, Corollary 1.4.3]. The proof in the real case requires a couple of additional observations. E.g. we need to add to Arveson’s proof the fact about real irreps from [29, Proposition 5.3.7 (1)]. Indeed the proof of [29, Proposition 5.3.7 (1) and Proposition 4.3.4 (3) and (4)] shows that $\varphi(V)^{\prime}$ is a real von Neumann algebra whose projections norm densely span its selfadjoint part. So $\varphi(V)$ is irreducible if and only there are no nontrivial such projections, or equivalently if $(\varphi(V)^{\prime})_{\rm sa}=\mathbb{R}\,I$. We also note that in Arveson’s proof(s) in [2, Section 1.4] we may assume that the operators $T_{i}$ there are selfadjoint, and where he says that $\pi(B)^{\prime}$ is scalar, we use $(\pi(B)^{\prime})_{\rm sa}$ is scalar. For ucp maps Arveson’s usage of ‘pure’ in [2] is equivalent to our Definition 2.2.
@@
For a $C^{*}$-algebra $B$ generated by a unital completely order isomorphic copy of $V$ we recall that the Shilov boundary ideal is the largest closed ideal $J$ of $B$ such that the quotient map $B\to B/J$ is completely isometric on $V$. (The quotient by the Shilov boundary ideal is a copy of the $C^{*}$-envelope.) Arveson showed in the complex case [2, Section 3] that the Shilov boundary ideal equals the intersection of the kernels of all boundary representations. The same proof works in the real case. His proof uses an ‘invariance principle’ which was simplified in [3, Proposition 3.1]. It also follows from e.g. the observation above [7, Proposition 4.1.12], which is the same in the real case. The proof of the following result from [13] is also the same in the real case.
@@
Let $V$ be a real operator system and $K$ its associated real nc convex set. Every real pure ucp map $\varphi:V\to B(H)$ can be dilated to an extreme ucp map $\widetilde{\varphi}:V\to B(L)$. Every real pure in $K$ has an extreme dilation.
@@
This is a combination of the real cases of Theorems 3.1, 3.3 and 3.4 in [13]. All of this (including the alternative arguments given there) works the same in the real case, with as usual $\mathbb{C}$ replaced by $\mathbb{R}$, and assuming all completely positive maps are selfadjoint. One argument there uses [23, Theorem 2.2], which states that for a real operator system $V$, finite $k$, and pure (selfadjoint) cp map $\psi:V\to M_{k}(\mathbb{R})$ there exists $m\leq k$, pure $\varphi\in{\rm UCP}(V,M_{m}(\mathbb{R}))$ and $\gamma\in M_{m,k}(\mathbb{R})$ with $\psi=\gamma^{\mathsf{T}}\varphi(\cdot)\gamma$. In turn this uses in the second paragraph of its proof some complex finite dimensional linear algebra, and some tricks that often fail for real spaces. Because of its independent interest we check this in the real case. Let $b=\psi(1)$ and $p$ the projection onto Ran $b$. To prove in that second paragraph that $p\in\psi(V)^{\prime}$ we proceed as follows. Consider completely positive $\psi_{c}:V_{c}\to M_{k}(\mathbb{C})$, and view $p,b\in M_{k}(\mathbb{R})\subset M_{k}(\mathbb{C})$. Indeed $p$ is then the projection in $M_{k}(\mathbb{C})$ onto $\mathbb{C}^{k}$. The proof in that second paragraph shows that $p\in\psi_{c}(V_{c})^{\prime}$, so that $p\in\psi(V)^{\prime}$. ∎
We also of course now obtain [14, Theorem 6.2.4] in the real case.
@@
Just as in the complex case, the $C^{*}$-algebra generated by the image of $V$ under a direct sum of sufficiently many extreme ucp maps (or boundary representations), is (a copy of) the $C^{*}$-envelope of $V$. Indeed this direct sum is maximal [14, Corollary 3.5]. Thus, a quick method to construct the real $C^{*}$-envelope of a real operator system $V$ is thus to take a maximal dilation $\rho$ of an inclusion $V\subset B(H)$. The $C^{*}$-algebra generated by $\rho(V)$ is the real $C^{*}$-envelope. (Note also that $\rho$ is maximal in $K_{c}$ by Lemma 2.7, so the $C^{*}$-algebra generated by $\rho_{c}(V_{c})=\rho(V)+i\rho(V)$ is a copy of the complex $C^{*}$-envelope of $V_{c}$.)

### v2

@@ Abstract.
We continue with the theory of real noncommutative (nc) convexity, following the recent and profound complex case developed by Davidson and Kennedy. The present paper focuses on the theory of nc extreme (and pure and maximal) points and the nc Choquet boundary in the real case, and on the theory of real nc convex and semicontinuous functions and real nc convex envelopes. Our main emphasis is on how these interact with complexification. For example some of our paper analyzes carefully how various notions of ‘extreme’ or ‘maximal’ interact with our earlier concept of the complexification of a convex set. Several new features appear in the real case, particularly in later sections of our paper, including the novel notion of the complexification of a nc convex function, and the complexification of the convex envelope of a nc function. With an Appendix by T. Russell.
2020 Mathematics Subject Classification:
Primary: 46A55, 46L07, 47A20, 47L07, 26B25; Secondary 46L51, 46L52, 47B92, 47L05, 47L25, 47L30,
@@ 1. Introduction
Very recently in [8] we developed the real case of the theory of real noncommutative (nc) convex sets in the sense of Davidson and Kennedy. We continue here with the theory of real noncommutative (nc) convexity, following the recent and profound complex case developed by those authors in the memoir [14]. As before, one of our primary goals is to develop the role and properties of the complexification in this context. Several motivations for our work are given in [8], one mentioned in more detail there is that ‘classical convexity’ is in some ways essentially a real theory, and so the real variant of the theory is likely to be important in the future in operator algebras, functional analysis, mathematical physics, and ‘quantum analysis’ (see e.g. [16] as a representative example of what part of this might look like). The present paper focuses on the real case of the topics from the middle sections of the deep memoir [14], namely on the theory of nc extreme (and pure and maximal) points and on Arveson’s nc Choquet boundary in the real case, and on the theory of real nc convex and semicontinuous functions and real nc convex envelopes; especially on how these interact with complexification. One use of this is as a tool to transfer results back and forth between the real and complex nc convexity theory, which often allows one to avoid difficulties. Some of our paper analyzes carefully how various notions of ‘extreme’ interact with our earlier concept of the complexification of a convex set. More sophisticated results, such as more advanced nc Choquet theory, will be discussed in a future paper. Several new features appear in the real case in later sections of our paper, including the novel notion of the complexification of a nc convex function on a nc convex set $K$, and the complexification of the convex envelope of a nc function on $K$. As just one example, we show that the convex envelope construction commutes with complexification. The convex envelope of a function $f$ defined on a convex set is of utmost importance in classical convexity theory, where it is sometimes called the ‘lower envelope’. (See e.g. [30]. Of course the upper/concave envelope may be treated via the convex envelope of f-f.) More generally we continue giving many foundational structural results for real nc convex sets and the functions on them.
@@
There is of course a large and highly impressive literature (going back thirty years or so) on noncommutative or quantum versions of extreme points, Krein-Milman, etc, and also on noncommutative versions of convex functions. As just a few recent examples see for instance the remarkable papers [20, 22, 34, 35] (and the list of papers referred to on p. 2 of [8]), and references therein. (At the time of ArXiV submission of the present paper Scott McCullough also pointed out the very very recent [21].) Indeed there are even classical precursors e.g. in the operator (or matrix) convex and monotone functions going back to Löwner. Moreover this large existing theory of ‘operator/matrix convexity’ of which we speak often focuses on real examples (e.g. spectrahedra), and sometimes is even completely cast in the real framework, using ‘real’ tools. However we believe that there is no technical overlap of our work with this literature, since our goal here is mainly to extend the perspective of [14] to the real case, particularly emphasizing and developing the role of the complexification in this endeavor. (In some stretches of Section 2 where the complexification plays little part we are essentially just verifying and resolving potential complications in, and otherwise expositing, the real case of [14, Section 6]. Our contributions in this part of Section 2 are not intended to be novel, but rather to be useful and systematic for future ‘real users’.) Indeed aspects of the work earlier cited in this paragraph will hopefully interact very profitably with the matter of the present paper(s) in the future, and in any case this cited work certainly furnishes very many interesting and important real examples. Every complex nc convex set or complex nc convex function is a real one, but there are many interesting and important real examples which are not complex ones. Also, much of the work cited earlier in this paragraph is stated in the Wittstock-Effros-Webster-Winkler language [40] of matrix convexity. We refer readers to [14] for discussion of some of the differences between nc and matrix convexity.
@@
Turning to the structure of our paper, in Section 2 we begin by characterizing maximal elements in a complexification. This gives a new route to the construction of the $C^{*}$-envelope of a real operator system (or unital operator space). We also discuss the real case of the theory of Davidson and Kennedy’s nc convexity variant of extreme, pure, and maximal points (which when ‘irreducible’ are essentially Arveson’s boundary representations). And of the associated nc Choquet and Shilov boundary, and its application to the nc Krein-Milman theorem and its converse. In most of the rest of this section we give a few points of clarification about the real case of the arguments, and note a few places where the arguments are different in the real case. In Section 3 we examine the connections between extreme, pure, and maximal points in the real and complex case, that is, how the these notions interact with the complexification of a nc convex set. One may view some of our work in this section as complementary to parts of the very recent paper [33]. In that paper some analogous results are established for $C^{*}$-extreme maps on real $C^{*}$-algebras. Every pure (and in particular any nc-extreme) ucp map on a $C^{*}$-algebra is $C^{*}$-extreme, but not vice versa. They do not treat general operator systems, thus their results would not apply to points in general nc convex sets. In Sections 4 and 5 we consider the real case of Davidson and Kennedy’s nc convex functions, including their multivalued nc convex functions, and convex envelopes. We give many foundational structural results here, for example developing the theory of the complexification in these settings and elucidating how it interacts with the basic objects and constructions. E.g. we show that the convex envelope construction commutes with complexification. As in [8], the complexification then allows us to derive the real variant of profound and extremely technical results from [14, Section 7] very quickly from the complex case.
@@
As usual, many of the results for real operator systems here or in [8] apply or have variants for the larger class of real unital operator spaces: If $V$ is a real operator space then $V+V^{*},V\cap V^{*}$, and $\{x+x^{*}:x\in V\}$ are well defined real operator systems [9, 10].
@@
We write $M_{n}(\mathbb{R})$ for the real $n\times n$ matrices, or sometimes simply $M_{n}$ when the context is clear. Similarly in the complex case. It is important that $n$ is usually not necessarily an integer, but rather a cardinal bounded by a fixed cardinal $\kappa$. We sometimes use the quaternions $\mathbb{H}$ as an example: this is simultaneously a real operator system, a real Hilbert space, and a real C-algebra, usually thought of as a real *-subalgebra of $M_{4}(\mathbb{R})$ or $M_{2}(\mathbb{C})$. Its complexification is $M_{2}(\mathbb{C})$. The letters $H,L$ are usually reserved for real or complex Hilbert spaces. Every complex Hilbert space $H$ is a real Hilbert space, i.e. we forget the complex structure. More generally we write $X_{r}$ for a complex Banach space regarded as a real Banach space. We write $X_{\rm sa}$ for the selfadjoint elements in a *-vector space $X$. In the complex case $M_{n}(X)_{\rm sa}\cong(M_{n})_{\rm sa}\otimes X_{\rm sa}$, but this fails for real spaces. A subspace of $B(H)$ is unital if contains the identity, and a map $T$ is unital if $T(1)=1$. Our identities $1$ always have norm $1$. For $z\in M_{n}(\mathbb{C})$ we write $\text{Re }z$ for $x\in M_{n}(\mathbb{R})$ where $z=x+iy$ for $y\in M_{n}(\mathbb{R})$. Finally, for a cardinal $n$ we define the isometry $u_{n}=\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$ where $1_{n}$ is the $n$-dimensional identity operator. We sometimes also write this as $u$.
The categories of real nc convex sets and real operator systems are equivalent (see [8, Section 4]), and we often switch without warning between such sets $K$ and the associated system, e.g. viewing $K=\mathrm{ncS}(V)$ for an appropriate real operator system $V$. A concrete complex (resp. real) operator system $V$ is a unital selfadjoint subspace of $B(H)$ for $H$ a complex (resp. real) Hilbert space. For $n\in\mathbb{N}$ we have the identification $M_{n}(B(H))\cong B(H^{(n)})$ where $H^{(n)}$ is the $n$-fold direct sum of $H$. From this identification, $M_{n}(V)$ inherits a norm and positive cone. The latter is the set $M_{n}(V)^{+}\coloneq\{x\in M_{n}(V):x=x^{*}\geq 0$ in $B(H^{(n)})\}$. For $n\in\mathbb{N}$ we define the amplification of a linear map $\varphi:V\to W$ by
@@
For a real operator space $E$, as in [14] we define $\mathcal{M}(E)=\bigsqcup_{n}\,M_{n}(E)$ (for $n$ cardinals bounded by some cardinal $\kappa$) with $M_{n}(E)$ the matrix space of $E$ for $X\subseteq\mathcal{M}(E)$. Here $\bigsqcup_{n}$ is the disjoint union. Define $X_{n}=X\bigcap M_{n}(E)$. In the case $E=\mathbb{R}$ write $\mathcal{M}=\mathcal{M}(\mathbb{R})$.
@@ 2. Extreme, pure, maximal points, and the nc Choquet boundary
Viewing $K=\mathrm{ncS}(V)$ for an appropriate real operator system $V$ as in [8, Section 4], this corresponds to $\psi\leq\varphi$ for a cp map $\psi$, implying that $\psi$ is a positive scalar multiple of $\varphi$. Indeed it is obvious that such a $\varphi$ is pure in the first sense above (note $x=t_{i}\alpha_{i}^{*}x_{i}\alpha_{i}$ evaluated at 1 implies $1=t_{i}\alpha_{i}^{*}\alpha_{i}$, so $\alpha_{i}$ is a multiple of an isometry). Conversely, if $\varphi=\varphi_{1}+\varphi_{2}$ for cp maps $\varphi_{i}$, use [9, Lemma 6.3] to write $\varphi=b_{1}\psi_{1}b_{1}+b_{2}\psi_{2}b_{2}$ for ucp maps $\psi_{i}$ and matrices $b_{i}\geq 0$ with $b_{1}^{2}+b_{2}^{2}=1$. If $\varphi$ is pure in the first sense then $\varphi_{i}=b_{i}\psi_{i}b_{i}$ is a positive multiple of $\varphi$. For most practical purposes this allows us in most situations to drop the requirement that $\alpha_{i}$ is a positive scalar multiple of an isometry above, instead saying that $\alpha_{i}^{\mathsf{T}}x_{i}\alpha_{i}$ is a positive scalar multiple of $x$. See Remark 6.1.6 in [14].
@@
A point $x\in K_{n}$ is called extreme if when it is written as a finite nc convex combination, $x=\sum_{i}\alpha_{i}^{\mathsf{T}}x_{i}\alpha_{i}$ for $x_{i}\in K_{n_{i}}$ and non-zero $\alpha_{i}\in M_{n_{i},n}(\mathbb{R})$ such that $\sum_{i}\alpha_{i}^{*}\alpha_{i}=1_{n}$, then $\alpha_{i}=c_{i}\beta_{i}$ where $\beta_{i}$ is an isometry and $c_{i}$ is a positive number such that $\beta_{i}^{\mathsf{T}}x_{i}\beta_{i}=x$ and $x_{i}$ commutes with $\beta_{i}\beta_{i}^{\mathsf{T}}$ for each $i$ (so that $x_{i}$ decomposes with respect to the range of $\alpha_{i}$ as $x_{i}=y_{i}\oplus z_{i}$ with $y_{i}$ unitarily equivalent to $x$). The set of real extreme points is denoted by $\partial K$ or Ch(K)(K).
We sometimes write the collection of extreme points as Ch(K)(K), since it corresponds to the Choquet boundary, not the Shilov boundary. The later Corollary 2.14 will say that these are essentially the same as the Arveson boundary representations.
Suppose that $\varphi:V\to B(H)$ is a unital complete order embedding on a real operator system $V$. The real $C^{*}$-algebra generated by the image of a maximal dilation $j:V\to B(L)$ of $\varphi$ is (a copy of) the important $C^{*}$-envelope $C^{*}_{\rm min}(V)$, by the well known arguments for this from [18, 3]. That is, it has the universal property that for any unital complete order embedding $\kappa:V\to B(L)$ there is a unique *-homomorphism $\pi:C^{*}(\kappa(V))\to C^{*}(j(V))$ with $\pi\circ\kappa=j$. The arguments (due to Dritschel-McCullough, Arveson, and others, see e.g. the historical notes to [7, Section 4.1]) for the next result work the same in both the real and complex case. As explained and proved in several places, e.g. in [3, Proposition 2.2] or [7, 4.1.11-4.1.13], we have:
@@
We next prove a useful characterization of the maximal elements of a complexification. We recall the fact from the first three lines of the last paragraph of the proof of [8, Theorem 3.4] that an element $x\in K_{2n}$ of a real nc convex set $K$ is of the form $c(a,b)$ for $a,b\in K_{n}$ if and only if for the unitary matrix $W=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$ we have $W^{*}xW=x$. In this case, $c(u^{*}xu)=x$. Real nc affine maps preserve such elements. Indeed if $\omega$ is nc affine then $W^{*}\omega(c(x))W=\omega(W^{*}c(x)W)=\omega(c(x))$. Hence also $c(u^{*}\omega(c(x))u)=\omega(c(x))$. We use this in the proof below.
@@
A direct proof of this result is possible (see e.g. the proof of Lemma 3.3). Instead we deduce it from swiftly from Theorem 2.6. By that result $x+i0\in K_{c}$ is complex maximal if and only if $x\oplus x$ is real maximal in $K$. However it is a pleasant exercise that this is equivalent to $x$ being maximal. ∎
@@
This gives a (new) route to the construction of the $C^{*}$-envelope of a real operator system (or unital operator space). Namely the real case of the Dritschel-McCullough-Arveson route (see [18, 3] or as mentioned before Theorem 2.4 or after Theorem 2.17), but using the ‘real arguments’ above.
@@
The discussion in 5.1.2 in [14] is the same in the real case. Turning to Section 6 of [14], Remark 6.1.3 in [14] and the following propositions have the same proof as in the complex case (see Proposition $6.1.4$ and $6.1.5$ in [14]).
@@
(1) This is essentially [2, Corollary 1.4.3]. The proof in the real case requires a couple of additional observations. E.g. we need to add to Arveson’s proof the fact about real irreps from [29, Proposition 5.3.7 (1)]. Indeed the proof of [29, Proposition 5.3.7 (1) and Proposition 4.3.4 (3) and (4)] shows that $\varphi(V)^{\prime}$ is a real von Neumann algebra whose projections norm densely span its selfadjoint part. So $\varphi(V)$ is irreducible if and only there are no nontrivial such projections, or equivalently if $(\varphi(V)^{\prime})_{\rm sa}=\mathbb{R}\,I$. We also note that in Arveson’s proof(s) in [2, Section 1.4] we may assume that the operators $T_{i}$ there are selfadjoint, and where he says that $\pi(B)^{\prime}$ is scalar, we use $(\pi(B)^{\prime})_{\rm sa}$ is scalar. For ucp maps Arveson’s usage of ‘pure’ in [2] is equivalent to Definition 2.2.
@@
For a $C^{*}$-algebra $B$ generated by a unital completely order isomorphic copy of $V$ we recall that the Shilov boundary ideal is largest closed ideal $J$ of $B$ such that the quotient map $B\to B/J$ is completely isometric on $V$. (The quotient by the Shilov boundary ideal is a copy of the $C^{*}$-envelope.) Arveson showed in the complex case [2, Section 3] that the Shilov boundary ideal equals the intersection of the kernels of all boundary representations. The same proof works in the real case. His proof uses an ‘invariance principle’ which was simplified in [3, Proposition 3.1]. It also follows from e.g. the observation above [7, Proposition 4.1.12], which is the same in the real case.
The proof of the following result from [13] is also the same in the real case.
@@
Let $V$ be a real operator system and $K$ its associated real nc convex set. Every real pure ucp map $\varphi:V\to B(H)$ can be dilated to an extreme ucp map $\widetilde{\varphi}:\mathcal{S}\to B(L)$. Every real pure in $K$ has an extreme dilation.
@@
This is a combination of the real cases of Theorems 3.1, 3.3 and 3.4 in [13]. All of this (including the alternative arguments given there) works the same in the real case, with as usual $\mathbb{C}$ replaced by $\mathbb{R}$, and recalling that all completely positive maps are selfadjoint. One argument there uses [23, Theorem 2.2], which states that for a real operator system $V$, finite $k$, and pure cp map $\psi:V\to M_{k}(\mathbb{R})$ there exists $m\leq k$, pure $\varphi\in{\rm UCP}(V,M_{m}(\mathbb{R}))$ and $\gamma\in M_{m,k}(\mathbb{R})$ with $\psi=\gamma^{\mathsf{T}}\varphi(\cdot)\gamma$. In turn this uses in the second paragraph of its proof some complex finite dimensional linear algebra, and some tricks that often fail for real spaces. Because of its independent interest we check this in the real case. Let $b=\psi(1)$ and $p$ the projection onto Ran $b$. To prove in that second paragraph that $p\in\psi(V)^{\prime}$ we proceed as follows. Consider completely positive $\psi_{c}:V_{c}\to M_{k}(\mathbb{C})$, and view $p,b\in M_{k}(\mathbb{R})\subset M_{k}(\mathbb{C})$. Indeed $p$ is then the projection in $M_{k}(\mathbb{C})$ onto $\mathbb{C}^{k}$. The proof in that second paragraph shows that $p\in\psi_{c}(V_{c})^{\prime}$, so that $p\in\psi(V)^{\prime}$. ∎
We also of course now immediately obtain [14, Theorem 6.2.4] in the real case.
@@
Just as in the complex case, the $C^{*}$-algebra generated by the image of $V$ under a direct sum of sufficiently many extreme ucp maps (or boundary representations), is (a copy of) the $C^{*}$-envelope of $V$. Indeed this direct sum is maximal [14, Corollary 3.5]. Thus, a quick method to construct the real $C^{*}$-envelope of a real operator system $V$ is thus to take a maximal dilation $\rho$ of an inclusion $V\subset B(H)$. The $C^{*}$-algebra generated by $\rho(V)$ is the real $C^{*}$-envelope. (Note also that $\rho$ is maximal in $K_{c}$ by Corollary 2.7, so the $C^{*}$-algebra generated by $\rho_{c}(V_{c})=\rho(V)+i\rho(V)$ is a copy of the complex $C^{*}$-envelope of $V_{c}$.)